1. The North American and European continents are moving apart at a rate of about 3 cm/y. At this how long will it take them to drift 800km farther apart than they are at present?

2. Assume air resistance is negligible unless otherwise stated. A coin is dropped from a hot-air balloon that is 250 m above the ground and rising at 12.0 m/s upward. For the coin, find the following.

a) the maximum height (in m) reached

b) its position and velocity (in m and m/s, respectively) 5.00 s after being released (Assume that the +x-axis is to the right and the +y-axis is up along the page. Indicate the direction with the sign of your answer.)

c) the time (in s) before it hits the ground
Answer #1

Only one question can be posted as per HomeworkLib guidelines.

1.

Given speed of continent moving

V=3 cm/y =0.03 m/y

Time taken by the continent to drift 800 km is

t =d/V =(800*103)/(0.03)=2.6667*107 years.

2.

a)

From Kinematic equation

Vfy2=Voy2-2g(y-yo)

At maximum height Vfy=0

0=122-2*9.8(Hmax-250)

Hmax=(144/19.6)+250 =257.3 m

b)

from

Y=Yo+Voyt-(1/2)gt2

at t=5 s

Y=250+12*5-(1/2)*9.8*52

Y=187.5 m

From

Vfy=Voy-gt

Vfy=12-9.8*5

Vfy=-37 m/s

c)

from

Y=Yo+Voyt-(1/2)gt2

when balloon hits the ground y=0

0=250+12t-(1/2)(9.8)t2

4.9t2-12t-250=0

t=8.47 s